OpenStudy (anonymous):

4√5/√30 four to the square root of 5 divided by the square root of 30

OpenStudy (anonymous):

2√6/3

sam (.sam.):

\[\frac{4\sqrt{5}}{\sqrt{30}}=\frac{4\sqrt5}{\sqrt5\sqrt6}=\frac{4}{\sqrt6}=\frac{4\sqrt6}{6}=\frac{2\sqrt6}{3}\]